# Generalized Bayesian Inference for Causal Effects using Covariate Balancing Procedure

### v2

@@ Appendix A Simulation Results
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
@@
Table A.5: Summary of estimated causal effects under the scenario (e): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
[…]
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 2.051 6.367 0.934 0.238 1.184 1.000 1.318 3.965 0.943 0.153 1.295 0.994 1.541 2.814 0.944 0.108 1.259 0.994
CBPS 1.664 6.409 0.933 0.238 0.961 1.006 0.995 3.996 0.943 0.154 0.977 1.002 1.213 2.835 0.944 0.109 0.991 1.001
RCAL CV 2.357 6.146 0.909 0.208 1.361 0.965 1.184 4.028 0.897 0.132 1.163 1.010 1.236 2.900 0.894 0.093 1.010 1.024
$\lambda=0.01$ 2.514 6.647 0.886 0.209 1.160 4.022 0.897 0.132 1.200 2.837 0.900 0.093
$\lambda=0.05$ 2.592 5.780 0.926 0.208 1.856 3.604 0.929 0.131 2.029 2.568 0.928 0.093
$\lambda=0.10$ 3.908 5.392 0.941 0.205 3.712 3.412 0.938 0.129 4.173 2.474 0.934 0.091
BRCAL PCIC 1.732 6.369 0.928 0.229 Ref. Ref. 1.018 3.987 0.932 0.147 Ref. Ref. 1.224 2.831 0.934 0.106 Ref. Ref.
$\omega=0.2$ 2.140 6.126 1.000 0.621 1.191 3.922 1.000 0.392 1.301 2.812 1.000 0.277
$\omega=0.5$ 1.834 6.290 0.999 0.393 1.067 3.966 0.998 0.248 1.247 2.826 0.999 0.175
$\omega=1.0$ 1.766 6.350 0.970 0.278 1.032 3.982 0.972 0.175 1.227 2.829 0.971 0.124
$\omega=1.5$ 1.730 6.369 0.926 0.227 1.016 3.987 0.927 0.143 1.226 2.832 0.925 0.101
Predictors Logit 7.883 8.079 0.916 0.269 1.337 1.028 4.404 5.067 0.934 0.182 1.173 1.028 3.593 3.566 0.942 0.131 1.239 1.018
CBPS 5.841 7.933 0.926 0.291 0.990 1.010 3.731 4.952 0.921 0.177 0.994 1.004 2.896 3.513 0.933 0.129 0.998 1.003
RCAL CV 7.678 7.404 0.854 0.208 1.302 0.942 5.147 5.324 0.820 0.132 1.371 1.080 4.384 3.964 0.788 0.093 1.511 1.131
$\lambda=0.01$ 21.105 19.361 0.722 372.870 7.462 6.823 0.741 0.133 4.864 4.211 0.760 0.093
$\lambda=0.05$ 9.754 9.576 0.802 0.339 3.206 4.024 0.904 0.132 2.347 2.800 0.907 0.094
$\lambda=0.10$ 5.735 5.976 0.915 0.208 4.067 3.472 0.939 0.130 3.832 2.485 0.934 0.092
BRCAL PCIC 5.898 7.858 0.855 0.230 Ref. Ref. 3.754 4.930 0.860 0.148 Ref. Ref. 2.901 3.504 0.860 0.105 Ref. Ref.
$\omega=0.2$ 6.123 7.423 1.000 0.619 3.808 4.786 1.000 0.392 2.955 3.451 1.000 0.277
$\omega=0.5$ 5.985 7.724 0.986 0.391 3.770 4.886 0.988 0.248 2.914 3.487 0.988 0.175
$\omega=1.0$ 5.912 7.830 0.922 0.276 3.755 4.920 0.925 0.175 2.909 3.500 0.923 0.124
$\omega=1.5$ 5.894 7.863 0.850 0.225 3.753 4.932 0.851 0.143 2.900 3.505 0.850 0.101
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method;
a: For CBPS, CIs were calculated from the results where errors did not occur. The number of errors was $\frac{2}{10000}(=0.02\%)$ when “Covariate” is Predictors.
Table A.6: Summary of estimated causal effects under the scenario (f): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 4.437 6.726 0.925 0.246 1.181 0.973 3.686 4.109 0.939 0.158 1.279 0.970 3.677 2.954 0.940 0.112 1.316 0.974
CBPS 3.680 6.953 0.923 0.254 0.980 1.006 2.846 4.247 0.941 0.165 0.988 1.003 2.782 3.038 0.942 0.117 0.995 1.001
RCAL CV 5.432 6.756 0.875 0.208 1.446 0.978 4.008 4.455 0.858 0.132 1.391 1.052 3.619 3.310 0.841 0.093 1.295 1.091
$\lambda=0.01$ 6.961 8.604 0.815 4.349 4.128 4.665 0.841 0.132 3.539 3.279 0.845 0.093
$\lambda=0.05$ 5.203 6.638 0.883 0.217 4.180 3.950 0.900 0.131 4.027 2.832 0.898 0.093
$\lambda=0.10$ 5.983 5.829 0.917 0.206 5.767 3.615 0.925 0.130 5.800 2.647 0.915 0.092
BRCAL PCIC 3.757 6.910 0.904 0.231 Ref. Ref. 2.882 4.236 0.915 0.148 Ref. Ref. 2.795 3.034 0.917 0.107 Ref. Ref.
$\omega=0.2$ 4.150 6.661 1.000 0.620 3.078 4.172 1.000 0.391 2.902 3.014 1.000 0.277
$\omega=0.5$ 3.869 6.832 0.997 0.392 2.937 4.214 0.997 0.247 2.819 3.026 0.996 0.175
$\omega=1.0$ 3.769 6.891 0.955 0.277 2.893 4.231 0.962 0.175 2.802 3.033 0.956 0.124
$\omega=1.5$ 3.750 6.911 0.897 0.226 2.876 4.236 0.905 0.143 2.794 3.034 0.903 0.101
Predictors Logit 7.883 8.079 0.916 0.269 1.337 1.028 7.660 4.960 0.936 0.179 1.342 1.010 8.304 3.612 0.929 0.129 1.308 1.018
CBPS 5.841 7.933 0.926 0.291 0.990 1.010 5.700 4.931 0.938 0.185 0.998 1.004 6.347 3.556 0.936 0.133 1.000 1.002
RCAL CV 7.678 7.404 0.854 0.208 1.302 0.942 11.113 6.732 0.733 0.136 1.947 1.371 13.423 5.528 2.114 1.558
$\lambda=0.01$ 21.105 19.361 0.722 372.870 16.553 9.576 0.642 0.161 15.519 6.210 0.618 0.100
$\lambda=0.05$ 9.754 9.576 0.802 0.339 6.723 4.842 0.829 0.132 6.724 3.388 0.834 0.093
$\lambda=0.10$ 5.735 5.976 0.915 0.208 5.170 3.695 0.920 0.131 5.655 2.709 0.912 0.093
BRCAL PCIC 5.898 7.858 0.855 0.230 Ref. Ref. 5.709 4.911 0.860 0.145 Ref. Ref. 6.350 3.549 0.860 0.104 Ref. Ref.
$\omega=0.2$ 6.123 7.423 1.000 0.619 5.732 4.784 1.000 0.391 6.376 3.506 1.000 0.276
$\omega=0.5$ 5.985 7.724 0.986 0.391 5.707 4.872 0.989 0.247 6.359 3.535 0.985 0.175
$\omega=1.0$ 5.912 7.830 0.922 0.276 5.708 4.902 0.924 0.175 6.351 3.546 0.919 0.123
$\omega=1.5$ 5.894 7.863 0.850 0.225 5.709 4.911 0.853 0.143 6.349 3.549 0.851 0.101
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Figure A.1: Summary of estimated causal effects: Bias, root mean squared error (RMSE), and 95% coverage probability (95%CP) in 10,000 iterations were summarized under “Predictor” situation. The left panel: $n=200$; the center panel: $n=500$; The right panel: $n=1000$. RCAL (CV) was used the CV described in Section 4.2. BRCAL (PCIC) was used the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Appendix B Proposed Method under some Covariate Missing
@@
$(Y_{1},Y_{0})\mathop{\perp\!\!\!\perp}(A,R)|\bm{X},\ \ \ R\mathop{\perp\!\!\!\perp}X_{1}|\bm{X}_{2}.$ (B.1)
@@
Therefore, from (B.1), a similar moment conditions in the main manuscript can be considered. For instance, we can consider the (2.2), replaced with following loss functions:

### v3

@@ B.3 Covariate balance diagnosis
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[4]; RCAL: Calibrating method proposed by Tan[10]; BLPS: Bayesian method proposed by Saarela et al.[Sa2015]; BRCAL: Proposed method;
(a): “Small complexity and good overlap,” (b): “Small complexity and poor overlap,” (c): “Moderate complexity and good overlap,” (d): “Moderate complexity and poor overlap,” (e): “Large complexity and good overlap,” (f): “Large complexity and poor overlap.”
B.4 Sensitivity Analysis for Simulation Experiments
Here, we conduct a sensitivity analysis (SA) for the simulation experiment results presented in the main manuscript. Specifically, we vary the hyperparameter for the prior distribution of $\lambda$. We examine three scenarios as follows. Note that all scenarios share the same prior mean of $0.1$, which is the common upper bound for the SMD.
Shape: 0.1, Rate: 1
Variance becomes: $0.1$; more strict setting compared to reference setting.
Shape: 0.01, Rate: 0.1
This setting is the same as the main simulation results in the main manuscript. Therefore, the expected value of simulation results in SA becomes similar as them. Variance becomes: $1$ (reference setting).
Shape: 0.002, Rate: 0.02
Variance becomes: $5$; more optimistic setting compared to reference setting.
The SA results are presented in Table B.5. In small sample sizes and simple propensity score model settings, a strict prior setting resulted in smaller bias; however, it also increased the variance of the posterior distribution, leading to wider confidence intervals (CIs). On the other hand, in other settings, such strict balancing does not appear to be necessary. Additionally, an optimistic prior setting does not introduce bias in the ATE estimates. In summary, we conclude that the effect of the prior setting for $\lambda$ on the ATE estimate is limited.
@@ B.4 Sensitivity Analysis for Simulation Experiments
Table B.5: Summary of estimated causal effects under the scenario (a)(f) (“Confounder” situation): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), and average length of CI (AvL) of the IPW estimator in 10,000 iterations were summarized for each scenario (in “Scenario” column). PCIC was used to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Scenario Shape Rate Bias RMSE 95%CP AvL Bias RMSE 95%CP AvL Bias RMSE 95%CP AvL
(a) 0.1 1 0.009 5.560 0.990 0.519 0.085 3.493 0.992 0.324 0.020 2.475 0.993 0.227
0.01 0.1 0.041 5.597 0.959 0.259 0.047 3.505 0.960 0.146 0.020 2.478 0.965 0.104
0.002 0.02 0.017 5.598 0.957 0.234 0.050 3.504 0.958 0.144 0.017 2.478 0.962 0.103
(b) 0.1 1 0.285 6.237 0.941 0.262 0.157 3.871 0.945 0.155 0.067 2.730 0.952 0.112
0.01 0.1 0.299 6.235 0.931 0.228 0.165 3.871 0.936 0.146 0.067 2.730 0.943 0.105
0.002 0.02 0.296 6.235 0.931 0.228 0.167 3.870 0.934 0.146 0.069 2.729 0.943 0.105
(c) 0.1 1 0.041 5.551 0.989 0.492 0.362 3.428 0.993 0.313 0.147 2.413 0.995 0.222
0.01 0.1 0.236 5.582 0.960 0.253 0.499 3.435 0.964 0.146 0.204 2.416 0.968 0.105
0.002 0.02 0.263 5.578 0.958 0.233 0.487 3.436 0.963 0.144 0.200 2.415 0.967 0.103
(d) 0.1 1 0.146 5.605 0.980 0.425 0.766 3.515 0.982 0.243 0.544 2.481 0.982 0.164
0.01 0.1 0.376 5.631 0.958 0.236 0.843 3.519 0.960 0.145 0.564 2.483 0.962 0.104
0.002 0.02 0.375 5.627 0.956 0.228 0.834 3.518 0.959 0.144 0.566 2.482 0.962 0.103
(e) 0.1 1 1.231 6.410 0.941 0.266 1.309 4.013 0.939 0.157 1.168 2.812 0.948 0.114
0.01 0.1 1.229 6.416 0.927 0.229 1.308 4.012 0.932 0.146 1.171 2.812 0.933 0.105
0.002 0.02 1.226 6.414 0.925 0.228 1.309 4.012 0.929 0.146 1.167 2.811 0.932 0.105
(f) 0.1 1 2.866 6.831 0.917 0.247 2.547 4.291 0.926 0.158 2.473 3.027 0.928 0.115
0.01 0.1 2.885 6.830 0.908 0.231 2.547 4.289 0.916 0.148 2.468 3.026 0.915 0.107
0.002 0.02 2.885 6.828 0.908 0.230 2.560 4.290 0.914 0.148 2.473 3.027 0.913 0.106
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[4]; RCAL: Calibrating method proposed by Tan[10]; BLPS: Bayesian method proposed by Saarela et al.[Sa2015]; BRCAL: Proposed method;
(a): “Small complexity and good overlap,” (b): “Small complexity and poor overlap,” (c): “Moderate complexity and good overlap,” (d): “Moderate complexity and poor overlap,” (e): “Large complexity and good overlap,” (f): “Large complexity and poor overlap.”
Appendix C Additional Materials for Real Data Analysis
In the real data analysis, we use the dataset known as the “Whitehall” dataset. Royston et al. [29] investigated the relationship between mortality and various factors, including socioeconomic status, using 10-year follow-up data from the Whitehall I study. This prospective cohort study included 18,403 male British civil servants working in London, excluding those in the Diplomatic Service and British Council. After excluding four participants with implausible systolic blood pressure readings and 106 with less than 10 years of follow-up, 17,260 men remained, of whom 1,670 (9.7%) died. Assuming non-informative censoring, logistic regression was used to estimate the odds ratio of death within 10 years in Royston et al. [29]. In their analysis, cigarette smoking and systolic blood pressure were treated as continuous risk factors, while age, plasma cholesterol concentration, and Civil Service grade were included as confounders.
Based on their manuscript, the same subjects are considered, with smoking status as the exposure variable and all-cause death within 10 years as the outcome. The baseline characteristics and standardized mean differences (SMDs) before and after confounding adjustment using inverse probability weighting (IPW) with ordinary logistic regression are summarized in Table C.6. As expected, the SMDs are smaller after adjustment than before. The absolute values of the SMDs are also plotted in Figure C.2. Combining these results with those in Section B.3, we expect that covariate balance is also achieved in our proposed method.
Table C.6: Demographic of Whitehall dataset
Non-smoking Smoking Overall SMD
$N=10103$ $N=7157$ $N=17260$ Before After
Age 51.3 (6.27) 52.0 (6.35) 51.6 (6.31) -0.116 -0.001
Cholesterol 5.11 (1.20) 5.10 (1.19) 5.11 (1.20) 0.007 -0.002
Job grade 1 682 (6.8%) 266 (3.7%) 948 (5.5%) 0.137 0.000
2 7539 (74.6%) 4478 (62.6%) 12017 (69.6%) 0.262 0.000
3 1265 (12.5%) 1447 (20.2%) 2712 (15.7%) -0.209 0.000
4 617 (6.1%) 966 (13.5%) 1583 (9.2%) -0.250 0.000
For continuous variables: mean (SD); for categorical variables: number of subjects (%);
Before: Before IP weighting; After: After IP weighting.
Figure C.2: SMD of confounders in Whitehall dataset.
Appendix D Proposed Method under some Covariate Missing
@@
$(Y_{1},Y_{0})\mathop{\perp\!\!\!\perp}(A,R)|\bm{X},\ \ \ R\mathop{\perp\!\!\!\perp}X_{1}|\bm{X}_{2}.$ (D.1)
@@
Therefore, from (D.1), a similar moment conditions in the main manuscript can be considered. For instance, we can consider the (2.2), replaced with following loss functions: